# Draw-In-Mind: Learning Precise Image Editing via Chain-of-Thought Imagination

### v1

@@ 4.2.2 Image Editing
| Gemini | 6.82 | 7.41 | 6.48 | 6.87 | 7.44 | 6.51 |
| GPT-4o | 7.40 | 7.90 | 7.14 | 7.22 | 7.89 | 6.98 |
| Doubao | 7.87 | 8.10 | 7.59 | 7.74 | 8.13 | 7.49 |
@@
| Instruct-P2P | 3.34 | 6.21 | 3.23 | 3.30 | 6.19 | 3.22 |
| MagicBrush | 4.56 | 6.34 | 4.24 | 4.52 | 6.37 | 4.19 |
| AnyEdit | 3.12 | 5.87 | 2.92 | 3.05 | 5.88 | 2.85 |
| OmniGen | 6.04 | 5.86 | 5.15 | 5.88 | 5.87 | 5.01 |
| UniWorld-V1 | - | - | - | 4.93 | 7.43 | 4.85 |
| Janus-4o* | 4.69 | 4.68 | 3.91 | 4.64 | 4.57 | 3.83 |
| Step1X-Edit | 7.29 | 6.96 | 6.62 | 7.13 | 7.00 | 6.44 |
@@
We further demonstrate the capability of DIM-4.6B-Edit through intuitive visual comparisons of editing results on four AI-generated out-of-domain images in Figure 3. Janus-4o exhibits severe distortions despite being trained on GPT-4o-generated edit pairs, while Step1X-Edit produces less natural edits (rows 2–4) and fails in complex scenarios such as row 1, which involves manipulating multiple objects. In contrast, DIM-4.6B-Edit successfully follows the instructions to produce natural and consistent edited images.
@@
We also include results on GEdit-Bench-EN in Table 5. Our model ranks just behind the in-domain tester, i.e., Step1X-Edit, while outperforming all other out-of-domain competitors. The task-wise performance, shown in Table 6, reveals a similar pattern to UniWorld-V1: Step1X-Edit achieves notable gains in the Text Change task, whereas other models, including ours, perform less effectively due to the absence of such data in DIM-Edit. Excluding the Text Change task, DIM-4.6B-Edit beats Step1X-Edit while maintaining a compact size, underscoring the high efficacy of our CoT data.
4.3 Ablation Study
| Stage1 Non-CoT Data | ||||||||||
| Stage2 CoT Data | ||||||||||
In Table 7, we present a rigorous data composition analysis for the editing task to identify the sources of performance improvements. In the first stage, we observe that training solely on ShareGPT-4o-Image already yields a satisfactory ImgEdit score, indicating strong semantic alignment, which is consistent with the behavior of Janus-4o. However, models trained exclusively on GPT-4o-generated data tend to alter the overall layout noticeably, which is undesirable. In contrast, training on UltraEdit produces slightly lower scores but preserves better consistency between the source and target images. When combining the two datasets, performance improves significantly, as the model benefits from the semantic richness while retaining the edit consistency.
The visualization of three variants finetuned from the base checkpoint ❀ in Table 7 is shown in Figure 3 for intuitive analysis. The variant tuned on ShareGPT-4o-Image (column 4) significantly alters the layout despite following the edit prompt, while its counterpart tuned on ShareGPT-4o-Image-CoT (column 5) preserves more details, indicating that CoT imagination helps maintain editing consistency. However, using ShareGPT-4o-Image-CoT alone still produces unstable edits. In contrast, the model tuned on the full DIM-Edit dataset (column 6), i.e., DIM-4.6B-Edit, achieves the best results in both semantic alignment and edit consistency, demonstrating the effectiveness of all three components in DIM-Edit.
| DIM-Edit | 4.09 | 3.47 | 2.30 | 4.00 | 3.43 | 3.87 | 4.92 | 2.85 | 4.08 | 3.67 |
| w/o GLP | 3.85 | 3.29 | 2.06 | 3.91 | 3.24 | 3.55 | 4.80 | 2.79 | 3.92 | 3.49 |
| w/o LOP | 3.80 | 3.15 | 1.92 | 3.83 | 3.07 | 3.60 | 4.79 | 2.44 | 3.92 | 3.39 |
| w/o EAL | 3.79 | 3.25 | 1.96 | 3.73 | 2.96 | 3.65 | 4.81 | 2.82 | 3.82 | 3.42 |
| w/o EII | 3.77 | 3.22 | 1.82 | 3.88 | 2.96 | 3.61 | 4.78 | 2.55 | 3.58 | 3.35 |
We also analyze the effect of each CoT component by individually removing it, with the results shown in Table 7. All components contribute positively to the final performance, though their importance varies. The Global Layout Perception (GLP) has only a minor impact, likely because it is an easy task for the generator. In contrast, the other three CoT components, i.e., Local Object Perception (LOP), Edit Area Localization (EAL), and Edited Image Imagination (EII), have a significant effect. LOP and EAL require the model to focus on specific regions, while EII demands complex reasoning; none of these are trivial for the generator. These findings further validate the Draw-In-Mind paradigm, which reduces the cognitive burden on the generator and thereby improves editing performance.
5 Conclusion
In this paper, we identify a crucial issue in existing image editing systems, i.e., imbalanced division of responsibilities, where the generator is burdened with complex reasoning, leading to reduced performance. To address this, we propose the Draw-In-Mind (DIM) dataset, consisting of two parts: (i) DIM-T2I, 14M web-crawled image–text pairs with carefully crafted long-context prompts that provide a foundation for complex CoT comprehension in editing; and (ii) DIM-Edit, 233K high-quality image editing pairs with detailed and precise CoT imagination. By training on the DIM dataset and incorporating an external designer during editing, we present DIM-4.6B-Edit, which achieves SOTA or competitive performance on ImgEdit and GEdit-Bench-EN while maintaining a tiny overall and trainable parameter size. These results validate our motivation to shift the design responsibility from the generation module to the understanding module, as well as the high efficiency of our proposed CoT edit data.

### v2

@@ Appendix A Additional Experiments
| Gemini (Comanici et al., 2025) | 6.82 | 7.41 | 6.48 | 6.87 | 7.44 | 6.51 |
| GPT-4o (OpenAI, 2025) | 7.40 | 7.90 | 7.14 | 7.22 | 7.89 | 6.98 |
| Doubao (Gong et al., 2025) | 7.87 | 8.10 | 7.59 | 7.74 | 8.13 | 7.49 |
@@
| Instruct-P2P (Brooks et al., 2023) | 3.34 | 6.21 | 3.23 | 3.30 | 6.19 | 3.22 |
| MagicBrush (Zhang et al., 2023) | 4.56 | 6.34 | 4.24 | 4.52 | 6.37 | 4.19 |
| AnyEdit (Yu et al., 2025) | 3.12 | 5.87 | 2.92 | 3.05 | 5.88 | 2.85 |
| OmniGen (Xiao et al., 2025) | 6.04 | 5.86 | 5.15 | 5.88 | 5.87 | 5.01 |
| UniWorld-V1 (Lin et al., 2025) | - | - | - | 4.93 | 7.43 | 4.85 |
| Janus-4o* (Chen et al., 2025b) | 4.69 | 4.68 | 3.91 | 4.64 | 4.57 | 3.83 |
| Step1X-Edit (Liu et al., 2025) | 7.29 | 6.96 | 6.62 | 7.13 | 7.00 | 6.44 |
@@
| Model | Gen. Resolution | Gen. Steps | Und. Params | Gen. Params | VAE Rate | Prompt | Speed |
| Step1X-Edit | 10241024 | 30 | 7B | 12.5B | 8 | Raw | 28.19s |
| DIM-4.6B-Edit | 3B | 1.6B | 32 | CoT | 6.23s |
Detailed Performance on GEdit-Bench-EN. Table 8 and 9 summarize overall and detailed task-wise performance of different models on GEdit-Bench-EN, respectively. Our DIM-4.6B-Edit ranks just behind the in-domain tester, i.e., Step1X-Edit, while surpassing all other out-of-domain competitors. Moreover, among out-of-domain testers, DIM-4.6B-Edit is the only model that consistently preserves both semantic consistency and perceptual quality. This demonstrates the effectiveness of DIM-Edit, where edits with high perceptual fidelity are precisely aligned with CoT-style imagination, thereby ensuring semantic correctness.
Inference Efficiency. Beyond precise image editing, our DIM-4.6B-Edit also maintains highly efficient inference inherited from the SANA architecture. To verify this, we compare the average editing time over 100 samples between Step1X-Edit and DIM-4.6B-Edit, as reported in Table 10. Specifically, Step1X-Edit is provided with short raw prompts, while DIM-4.6B-Edit is evaluated with longer CoT prompts. Even under this more demanding setting, our model achieves a 4.5 speedup while preserving high editing quality, highlighting the effectiveness of the proposed DIM dataset and the Draw-In-Mind paradigm.
Visualizations of Different Editing Operations. Beyond Figure 3 in the manuscript, we further visualize the edits of Janus-4o, Step1X-Edit, and our DIM-4.6B-Edit under the operations of add, change, remove, replace, and style transfer in Figure 4, 5, 6, 7, and 8, respectively. As shown, DIM-4.6B-Edit consistently preserves the overall layout while performing natural edits. For instance, in Figure 4, Janus-4o fails to generate details of the wooden cabin, while Step1X-Edit places the chimney on the river, which is counterfactual. In contrast, our DIM-4.6B-Edit carefully adds the wooden cabin while ensuring naturalness. In Figure 5, Janus-4o fails to follow the color change instruction. Step1X-Edit changes the singer’s shirt to blue but also alters fine details such as the collar shape. By comparison, our DIM-4.6B-Edit changes the shirt to red while preserving all details, including the shadow cast by the hand. In Figure 6, both DIM-4.6B-Edit and Step1X-Edit perform successful removals, whereas Janus-4o fails to remove the seaplane. In Figure 7, only DIM-4.6B-Edit captures the semantics of “majestically” and generates a roaring lion. Finally, in Figure 8, although all three models succeed in style transfer, only DIM-4.6B-Edit captures subtle visual cues, such as the green grass in the last row, and repaints them faithfully in the edits.
Appendix B DIM-T2I Analysis Dimensions
Figure 9 and 10 illustrate the 21 analysis dimensions and their corresponding prompts used in the DIM-T2I annotation process. These dimensions were carefully selected to encompass both broad world knowledge and fine-grained visual attributes, such as illumination and shadow effects, which are crucial for high-quality image generation. As a result, DIM-T2I provides not only semantically rich supervision but also consistent aesthetic guidance, forming a strong foundation for precise image editing. This is further demonstrated by the trained DIM-4.6B-T2I model, which faithfully follows user instructions at the semantic level while producing perceptually high-quality images under limited parameter budgets, achieving high GenEval scores and low FID on MJHQ-30K.
Appendix C DIM-Edit Data Collection Pipeline
As stated in Section 3.1.2, we collect raw edit data from four publicly available datasets:
UltraEdit (Zhao et al., 2024). To ensure editing consistency, we retain only those edit pairs that satisfy the following conditions: (i) the CLIP similarity between the source and edited images is greater than 0.9; (ii) the DINOv2 similarity is greater than 0.9; (iii) the SSIM score is greater than 0.8; and (iv) the prompt does not contain “rainbow”, since many edit pairs meeting (i)–(iii) are associated with low-quality “rainbow” edits. After filtering, we obtain roughly 160K edit pairs.
MagicBrush (Zhang et al., 2023). We include only 8K images from the training set to avoid potential information leakage during evaluation.
SEED-Data-Edit-Part3 (Ge et al., 2024). Since the “remove” operation is absent in UltraEdit, we additionally select 19K edit pairs from SEED-Data-Edit-Part3 by filtering prompts that explicitly contain “remove.”
ShareGPT-4o-Image (Chen et al., 2025b). We include only its 46K image-to-image subset.
By combining these collected datasets, we obtain a total of 233K raw edit pairs for the proposed DIM-Edit.
Appendix D DIM-Edit Quality Assessment
We further assess the quality of the CoTs in DIM-Edit through MLLM-powered validation. Specifically, due to API quota limitations, we randomly sample 30K edit pairs from DIM-Edit and use GPT-4.1 to evaluate the quality of the GPT-4o-annotated CoTs, categorizing them into four levels:
Low: The optimized edit instruction does not reflect the change between the source and edited images at all.
Medium: The optimized edit instruction captures the major change between the source and edited images, but the chain-of-thought contains some factual errors.
High: The optimized edit instruction captures the major change between the source and edited images, and the chain-of-thought contains only minor factual errors.
Ultra-High: The optimized edit instruction accurately captures all changes between the source and edited images, and the chain-of-thought contains no factual errors.
The percentage distribution of each quality level is shown in Figure 11. Notably, no data is categorized as “Low”, while the majority falls under the “Ultra-High” level, demonstrating the strong overall quality of DIM-Edit.